OpenStudy (anonymous):

prove that determinant is equal to (a-1)^3(a+3),where R1=[a 1 1 1],R2=[1 a 1 1],R3=[1 1 a 1],R4=[1 1 1 a].

OpenStudy (anonymous):

plz solve my question!

OpenStudy (anonymous):

Laplace's formula :D

OpenStudy (anonymous):

can you solve it by column or row operation?

OpenStudy (anonymous):

I'm doing it as follows: first I found determinant of (b 1 1) (1 a 1) (1 1 a) and got (a-1)(ab+b-2). Then your determinant is equal to a det (a 1 1) - 1 det (1 1 1) + 1 det (1 a 1) - 1 det( 1 a 1) (1 a 1) (1 a 1) (1 1 1) (1 1 a) (1 1 a) (1 1 a) (1 1 a) (1 1 1) First determinant is (a-1)(a^2+a-2) Second determinant is (a-1)(a+1-2) = (a-1)^2 Third determinant is -(a-1)^2 (swapping two rows changes sign of determinant) Fourth determinant is (a-1)^2 (same reason) Hence answer is a(a-1)(a^2+a-2) - (a-1)^2 - (a-1)^2 - (a-1)^2 = a(a-1)(a^2+a-2) - 3(a-1)^2 = (a-1)[a(a^2+a-2) - 3(a-1)] = (a-1)[a^3 + a^2 - 2a - 3a + 3] = (a-1)[a^3 + a^2 - 5a + 3] = (a-1)(a+3)(a^2-2a+1) using long division = (a-1)(a+3)(a-1)^2 = (a-1)^3(a+3) as required.